OpenStudy (anonymous):

Solve logx + 5 36 = 2

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Answer is 1

Directrix (directrix):

x = 6 - 5 = 1. The value x = -6 -5 = -11 does not apply given that log bases cannot be negative.